OpenStudy (anonymous):

If f(x) = 2(x)^2 + 5 square root of (x+2)

hartnn (hartnn):

whats the entire Question ?

OpenStudy (anonymous):

Completed the following statement : f(0) =?

hartnn (hartnn):

to find f(0), just put x =0 in f(x) so, \( \Large f(0) = 2(0)^2 + 5 \sqrt{ (0+2)} = ..... ?\) can you find ?

OpenStudy (anonymous):

I did that but my answer isn't coming out right.

hartnn (hartnn):

did u get \(\Large 5\sqrt2\) as your answer ?

OpenStudy (anonymous):

Yeah doi not simplify that?

hartnn (hartnn):

thats the most simplified form :)

OpenStudy (anonymous):

So that's what f(0) equals,

hartnn (hartnn):

yup. thats it!

OpenStudy (anonymous):

I'm doing a test online and it doesn't let me do square roots, so I don't think that's the answer they want.

hartnn (hartnn):

firstly, it is not allowed to ask test/graded questions on OpenStudy, refer to code of conduct, http://openstudy.com/code-of-conduct. secondly, if your function is f(x) = 2(x)^2 + 5 square root of (x+2) then f(0) surely is 5 square root 2, and your system should allow that answer.

OpenStudy (anonymous):

It's a practice quiz. It's open note. And you can ask for help. Sorry shouldn't of said test.

hartnn (hartnn):

then it is ok. anyways, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

Thanks for your help

hartnn (hartnn):

you're welcome ^_^